OpenStudy (anonymous):

Simplify. x/4x + x^2 A. 1/4 B. 1/x + 4; where x / -4, 0 C. 1/4 + 1/x; where x = 0 D. 1/x; where x = 0 Thank you.

OpenStudy (anonymous):

Oh and all of the = signs have a little line through them.

OpenStudy (anonymous):



OpenStudy (campbell_st):

factorise the denominator \[\frac{x}{x(4 + x)}\] remove the common factor of x now you can't have a zero denominator so x cannot equal 0 or 4 as this makes the expression undefined

OpenStudy (campbell_st):

oops should have written it as \[\frac{x}{x(x +4)}\]

OpenStudy (anonymous):

So is it C?

OpenStudy (campbell_st):

oops another typo you can't have 0 or -4 in the denominator the fraction becomes \[\frac{1}{x +4}\]

OpenStudy (campbell_st):

choose B

OpenStudy (anonymous):

Oh okay. I almost said B but you said 4 before. Thanks for pointing that out to me. Thanks for the help.